OpenStudy (anonymous):

A cable installer charges $40.00 per hour plus a $60.00 service charge. Find the total charges if it takes the cable installer 8.5 hours to complete the task

OpenStudy (anonymous):

40h+65=? 40(8.5)+65=? 340+65=? $405=? Answer is $405.00

OpenStudy (anonymous):

Thanks